Problem 1. Express the function 100-100n2+100n+3 in terms of e notation. Completely justify your answer
Answer #1
f(n) = Θ (g(n)) means there are positive constants c1, c2, and k, such that 0 ≤ c1g(n) ≤ f(n) ≤ c2g(n) for all n ≥ n0.
The values of c1, c2, and k must be fixed for the function f and must not depend on n.

(n^3 / 1000) - 100n^2 + 100n + 3 = Θ(n^3)

1)
(n^3 / 1000) - 100n^2 + 100n + 3 <= c(n^3)
Let's assume c = 2/1000
(n^3 / 1000) - 100n^2 + 100n + 3 <= 2/1000(n^3)
- 100n^2 + 100n + 3 <= 1/1000(n^3)
- 100n^2 + 100n + 3 <= 1/1000(n^3) is true for all n >= 0

2)
(n^3 / 1000) - 100n^2 + 100n + 3 >= c(n^3)
Let's assume c = 1/2000
(n^3 / 1000) - 100n^2 + 100n + 3 >= 1/2000(n^3)
(n^3 / 2000) - 100n^2 + 100n + 3 >= 0
(n^3 / 2000) - 100n^2 + 100n + 3 >= 0 is true for some positive value

so, (n^3 / 1000) - 100n^2 + 100n + 3 = Θ(n^3)

Answer: Θ(723)